Q. A hemisphere and a solid cone have a common base.The centre of mass of the common structure coincides with the centre of the common base.If R is the radius of hemisphere and h is the height of the cone,then


what is the value of h/R ?

Grade:11

4 Answers

Aman Bansal
13 years ago

dear raj,

com of cone= h/4 from base

and com of hemisphere = 5r/8 from base

since com of both of them coincides therefore

h/r =  5/2


We are all IITians and here to help you in your IIT JEE preparation.
Now you can win

exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.
Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar :
Click here to download the toolbar

 

AMAN BANSAL

Ankita
5 years ago
😃..Was it real Aman Bansal #sir..
Topper of JEE Advance 2016??
7 years ago🙄😓😧😧😧😧 he was in8th /9th std. COM kese padha..nahi....?😵😵
 
Ankita
5 years ago
7 saal pehle kese IITian ban gae?.
He might be someone else 🙃.
Got muddle headed.
Its ok............
Ankita
5 years ago
By the way the answer is √3..
I do know a bit bout bhe solution but not exactly .whatever ...the answercs √3... manipulate it yourself😃😆